OpenStudy (lgbasallote):

can someone give me some *possible to solve* integrals? i'd like to practice :)

OpenStudy (lgbasallote):

and please no mimintegrals or omnintegrals :p

OpenStudy (anonymous):

\[\int\limits \sqrt{a^{2}-x^{2}}dx\] i suppose u can do that!!

OpenStudy (mimi_x3):

Here's a fun one: \[\int\limits\frac{1}{\ln(x+1)} dx\] Or: \[\int\limits\sqrt{tanx} dx\] (without partial fracrtions)

OpenStudy (anonymous):

\[\int\limits_{1}^{2}\ln x/x ^{2}\]

OpenStudy (mimi_x3):

No mimintegrals? :(

OpenStudy (lgbasallote):

i said no mimintegrals or omnintegrals @Mimi_x3 :p

OpenStudy (mimi_x3):

I feel discrimininated :P

OpenStudy (mimi_x3):

I shall insert a complaint for the discrimination :P

OpenStudy (lgbasallote):

\[\int \frac{\ln x}{x^2}dx\] u = lnx du = 1/x dx dv = 1/x^2 dx v = -2/x^3 dx \[\frac{-2\ln x}{x^3} - \int -\frac{2}{x^4}dx\] it's integrable now

OpenStudy (lgbasallote):

\[\int \sqrt{a^2 - x^2} dx \implies \sin (\frac{x}{a})\] right?

OpenStudy (lgbasallote):

i mean \[\sin^{-1} (\frac{x}{a})\]

OpenStudy (anonymous):

nope!!

OpenStudy (mimi_x3):

My integral is left isolated..

OpenStudy (lgbasallote):

\[\int \frac{1}{\ln (x+1)}dx\] let u = x +1 du = dx \[\int \frac{du}{\ln u}\] you do the rest :p

OpenStudy (mimi_x3):

Wrong!

OpenStudy (lgbasallote):

heh :p

OpenStudy (mimi_x3):

Not thateasy (:

OpenStudy (lgbasallote):

that's why i left it haha

OpenStudy (mimi_x3):

However, the second one \[\int\limits\sqrt{tanx} dx\] without partial fractions is doable! (:

OpenStudy (lgbasallote):

lol :p

OpenStudy (mimi_x3):

Atleast try! you wanted practice.

OpenStudy (lgbasallote):

tell me how to get the integral of sin^4 x dx then ill give you that integral in a heartbeat :p

OpenStudy (mimi_x3):

Reduction formulae..

OpenStudy (mimi_x3):

http://en.wikipedia.org/wiki/Integration_by_reduction_formulae I gotta review these lol

OpenStudy (lgbasallote):

is there no one else?

OpenStudy (anonymous):

\[\large\int\limits (x^{3m}+x^{2m}+x^m) (2x^{2m}+3x^m+6)^{1/m} dx\]

OpenStudy (lgbasallote):

uhhh nevermind im closing this now :S lol

OpenStudy (lgbasallote):

you guys have already proven you are capable of giving hard ones :P lol

OpenStudy (anonymous):

that is a nice one let u=x^m

OpenStudy (anonymous):

@mukushla .....is is an awsome one!!!!!

OpenStudy (anonymous):

@lgbasallote \[(\sin^{2}x+\cos^{2}x)^{2}=1\] i think this is the starting for ur problm